I need help finding the answer to this question.
The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs.

Answer by Paul(988)

On le g= 4x-8
EQUATION:
2(4x-8)+x=29
8X-16+x=29
9x=45
x=6
4(6)-8=16
Hence, the sides are 16cm each and the base is 6cm,
Paul.